What is the pH of a 0.030 M solution of weak acid HA for which Ka = 6.6 x 10-5?
Answer #1

A weak acid is partially dissociated is water:

AH \rightleftharpoons A^{-} + H^{+}

The pH of a weak acid can be caclculated (based on Mass Action Law: K_{a}=\frac{[H^{+}]\times [A^{-}]}{[A]} ) using the following expression:

pH =\frac{1}{2}\left ( pK_{a} - logC\right )= \frac{1}{2}\left (-log(6.6\times 10^{-5}) - log0.03\right )=3.35